# MLLM-Tool: A Multimodal Large Language Model For Tool Agent Learning

### v2

@@ 3 Dataset Construction
3.1 API Collection
Initially, we crawled 263,945 models from the HuggingFace platform. However, the quality of these models varies significantly, for example, some models are poorly maintained, even lacking information in the model card. To ensure only those models meeting specific criteria (e.g. having a detailed model card, especially providing model description and example script) are ultimately selected, we employ a meticulous filtering process and the detailed filtering process is shown in the Appendix. Finally, we only retain 932 high-quality models.
For each model, we document the model card and extract the key information, including the domain, api__\__name, api__\__call, parameters, coarse__\__functionality, fine__\__functionality, example codes, and descriptions. Further, to prevent an overabundance of APIs with identical functionalities leading to excessive options for subsequent instruction matching (e.g., in the imaging realm, where two APIs might differ only in their network architectures but are identical in terms of the datasets they were trained on and their functionalities), we adopt a strategy based on the API’s download statistics. For any single functionality category, we retain a maximum of five of the most downloaded APIs.
It is worth noticing that the function classification rules formulated by the HuggingFace platform are coarse-grained and not suitable for direct use by our work for some reasons. In particular, some models lack assigned categories or are incorrectly classified, or they contain multiple tasks but only one of them is labeled. Additionally, the platform employs overly coarse, resulting in a loss of fine-grained distinctions. For example, the task of Audio Classification, can be subdivided into Event Recognition, Command Recognition, Spoken Language Identification, Speaker Verification, etc. We then propose a new hierarchical function categorization system and rearrange the filtered model. We list all the tasks our system supports in Table 2, and for the subtasks we divide under these tasks, we provide more detailed information in the Appendix.
Audio Classification Audio-to-Text Audio-to-Audio
Voice Activity Detection Video Classification Video-to-Text
Depth Estimation Image Classification Image Segmentation
Image-to-Image Visual Question Answering Image-to-Text
Zero-Shot Image Classification Multimodal-to-Image Object Detection
Feature Extraction Fill-Mask Text Generation
Text Generation Text Classification Question Answering
Sentence Similarity Summarization Text-to-Image
Text-to-Speech Text-to-Video Text-to-Text
Translation Zero-Shot Classification
Table 2: The 29 coarse-grained tasks our system supports.
Given our API collection may not cover all tasks, users in real-world scenarios might encounter situations without a suitable API. Thus, we also introduce the “Unknown” option to address this issue.
3.2 Instruction-Answer Pairs Generation
After meticulously filtering and refining through manual review, 10 well-crafted queries will be finally selected to be associated with each API call. A significant highlight of our approach is its capability to disambiguate text prompts by learning the characteristics of other modalities when handling multi-modal inputs. We elaborate on five text ambiguity cases that need to be checked:
Data models used are from different domains: It is imperative to acknowledge the inherent disparities across data domains, either auditory or visual. Owing to these pronounced differences between domains, a unified treatment could result in significant biases in outcomes. Models need to judiciously select APIs contingent on the particular data domain at hand, instead of using a unified treatment, ignoring the bias across different domains. For instance, in image segmentation tasks, it is vital to select distinct APIs tailored for specific scenarios like indoor decorations, portraits’ fashion apparel, or medical related according to the given image.
Image Quality: The most evident instance revolves around image quality, which can be compromised by multiple factors including fog, rain, low resolution, and motion blur. When confronted with vague commands such as “I want to enhance the image quality”, analyzing the reason for the low imaging quality of the input image, and then selecting the appropriate API call for quality enhancement becomes necessary.
Input Conditions of different models: In some cases, the format of the input image can also assist in resolving text ambiguity issues. For example, ControlNet  [45] posits multiple conditions for image generation. Faced with an instruction like “Based on the provided image, generate a photo of a chef in the kitchen,” and given inputs such as a pose image or a normal map, the model must leverage the unique features distinguishing between the two types of images. This discernment facilitates the selection of the appropriate API to achieve the desired output.
Figure 3: The word cloud calculated based on the frequency of task name occurrence.
Others: There are some individual cases unique to individual modality, for instance, linguistic variations in the audio realm. Within the speech recognition task, when presented with an instruction like “In my language class, we listened to a story. I managed to record it; can you tell me its content?”, the absence of explicitly provided language information necessitates the model to first ascertain the spoken language. Only then can it aptly select the corresponding API for accurate content extraction and interpretation.
We collect a total of multi-modal instruction-answer pairs, including 3798 image inputs, 103 video inputs, and 726 audio inputs, totaling 72 videos, 3777 images, and 590 audio files.
3.3 Instruction Matching
In contrast to prior works, our dataset stands out with a unique feature: it encompasses numerous API calls that perform identical or similar functions. This results in a scenario where a single instruction could be associated with several API calls within our dataset, each potentially suitable for addressing it. In step two, we approach from a singular API function perspective to construct one-to-one matched instruction-answer pairs, now necessitating integration of results from the classification system established in the first step. For APIs with identical functionalities, we can match their queries immediately. However, for groups of APIs with similar functionalities, it’s essential to determine whether a query aligns with a strict or broad functional description of each API. For instance, the request, “Can you help me generate an image of a ’golden-haired magical princess’ in the Disney style?” In this scenario, the emphasis on the “Disney style” narrows down the choice of APIs to those pre-trained on Disney-themed images. Conversely, a more general instruction like, “I’m organizing a role-playing event and need a reference image. Can you generate an image of a ’masked mysterious hero’ for me?” leaves room for multiple APIs to be potential solutions, given its lack of specificity. This process yielded a dataset where each instruction corresponds to one or, in some cases, multiple relevant API calls.
Figure 4: Our network architecture. The network could processes four distinct types of input, each uniquely color-coded. These inputs, irrespective of their modality, are first encoded using modality-specific frozen encoders. Subsequently, they are passed through a trainable projection layer, aligning them within a unified language feature space. The fine-tuned Large Language Models (LLMs), augmented with Low-Rank Adaptation (LoRA), then process a combination of the prompt and the projected embeddings, enabling the accurate predicting of corresponding API name.
3.4 Dataset Statistics
In alignment with HuggingFace’s task-specific API distributions, ToolMMBench exhibits a similar long-tailed distribution, with text-related APIs being the most abundant and video-related ones the least. ToolMMBench consists of 27 coarse-grained task categories inheriting from HuggingFace, from which we expands to 29 tasks to form the first level of our classification system, and then extending to our predefined fine-grained tasks. The overall distribution graph of our support tasks can be observed in Figure 3, where the larger the font size, the higher the frequency of occurrence. We also tally the instances of each ambiguity type in Figure 1(b), amassing a total of 2533 ambiguity pairs. Predominantly, these are concentrated within conditional cases involving image inputs, comprising 1530 pairs, since our dataset covers 11 distinct condition types of ControlNet  [45], and their various combinations, achieving a considerable quantity.
We divide our dataset into training and testing sets in an 8:2 ratio. Since our task currently focuses on a closed set of APIs, our test set cannot include any APIs that haven’t been encountered in the training set. Therefore, we approach this from the perspective of queries generated by each API. Given the presence of one-to-many query scenarios in our dataset, we implement a strategy in the training set to address these cases. Specifically, we split each one-to-many query into W distinct inputs, where W represents the number of APIs that support the given query. This approach ensures that each potential API response is individually represented, allowing the model to learn the diverse range of suitable responses for a single query. In contrast to the training sampling strategy, for the one-to-many query scenarios in our dataset, each query appears only once in the test set. Finally, the training set comprises 44,533 instructions, containing 9,245 different instructions, and the testing set totaled 2,397 instructions. The detailed distribution is illustrated in Table 3.
Settings Text Audio Image Video
Training Set (w/ split) 31,280 2,944 10,085 224
Training Set (w/o split) 5,563 562 3,938 82
Testing Set 1,452 164 760 21
Table 3: Distribution of training set and testing set. Note that, in the training set, one-to-many scenarios are split into multiple one-to-one instructions, so the first line represents for the number of queries after splitting while the second line calculates the distinct query number of four single modality. The last line is the distribution of the whole testing set.
@@ 3.4 Dataset Statistics
Particularly for the testing set, we depict in Figure 1(a) and Figure 1(c) the distribution of API quantities across each testing subset. These subsets are constructed based on three distinct partitioning criteria: ambiguity types, API option number, and modality.
4 MLLM-Tool
We introduce MLLM-Tool, a method that leverages the LLM performing the role of multi-modality tool agent. The architecture of this network is depicted in Figure 4. Considering the diversity of our data, which includes images, videos, and audios, we incorporate ImageBind  [9] as our primary multimodal encoder. ImageBind  [9] is designed to create a unified embedding space that integrates six distinct modalities, utilizing the advancements in large-scale vision-language models.
In our dataset. there are four types of modality-combination input: Text, Text+Image, Text+Video, and Text+Audio. The output from the LLM includes textual responses that provide tool recommendations along with their respective model card information. For the multimodal inputs in the Image, Video, or Audio modalities, we first utilize ImageBind  [9] encoders with fixed weights for feature extraction. These features are then aligned with the LLM’s feature space through a linear projection layer. To optimize our model while minimizing the number of learnable parameters, we implement Low-Rank Adaptation (LoRA)  [11] to fine-tune an open-source LLM.
Currently, the system’s output only includes the API Name in the format {api_name: [Recommended API]}. Our future endeavors aim to expand this output to include API parameters, code execution results, and elements of Chain-of-Thought (CoT) Prompting  [35]. Additionally, the current version of our system is designed for single-turn dialogues, with the intention to extend its capabilities to multi-turn conversations in subsequent developments.
5 Evaluation Metrics
  • Ambiguity types: In light of the five distinct ambiguity scenarios delineated above, we construct corresponding testing subsets. These facilitate the evaluation of the model’s capability to leverage multimodal information under varying types of ambiguity, providing a nuanced understanding of its performance.
  • Multiple options: Since our model has the capacity to support multiple options, we divide the instruction set into two subsets, solved only by one single API versus those allowing for multiple API options, and we expect to examine the performance gaps within two subsets.
  • Individual modality: For each modality, we partition subsets to evaluate model performance with fixed modality combinations, ensuring a comprehensive assessment across varying input configurations.
The testing set is divided into several subsets based on the three attributes. For a subset $\mathbf{S}$, supposing containing $n$ instructions $\{x_{1},x_{2},x_{3}\dots x_{n}\}$, where each instruction $x_{i}$ matches an API list $y_{i}$, which consists of a list of $M$ suitable APIs $[A_{i}^{1},A_{i}^{2},A_{i}^{3},\dots,A_{i}^{M}]$. Thus, the collection of API lists corresponding to this instruction subset combined to a set $\mathbf{Y}$, $\{y_{1},y_{2},y_{3}\dots y_{n}\}$. With these symbols defined, assuming the subset $\mathbf{S}$ yields an API set $\{z_{1},z_{2},z_{3}\dots z_{n}\}$, output from our model, we can represent the accuracy of each mentioned subset using the following formula:
$\text{Acc}=\sum\limits_{i=1}^{n}\mathbbm{1}_{z_{i}\in y_{i}}/n$ (1)
The above formula also applies to the accuracy calculation for the entire testing set. Besides that, the hallucination rate is used as a metric to evaluate the performance of each testing subset and the overall testing set. For the calculation of the hallucination rate, we first assume the corpus C, containing all APIs. Hallucination refers to the phenomenon where the model outputs APIs are not present in the corpus C, essentially fabrications of the model. Taking the above subset $\mathbf{S}$ as an example, the calculation of the hallucination rate can be expressed with a formula:
$\text{Hallu}=\sum\limits_{i=1}^{n}\mathbbm{1}_{z_{i}\notin C}/n$ (2)
Besides that, unlike simple one-to-one mapping relationships, using single-instance accuracy is insufficient for evaluating the system’s comprehension of the relationships of APIs, especially in one-to-many mapping relationships. Hence, we propose utilizing recall, allowing the system to perform repeated inferences. By examining the proportion of ground truth encompassed by the collective results of multiple inference results, marked as $R_{i}^{j}$, where $i$ is the $i^{th}$ instruction in the testing set and $j$ represents for the $j^{th}$ inference. The corresponding metric calculation is expressed as follows:
$\small\text{Recall}=\frac{1}{n}\sum_{i=1}^{n}\frac{(f_{c}(Set({R_{i}^{1},R_{i}^{2}\dots R_{i}^{K}|R_{i}^{j}\in y_{i},j\in[1,K]}))}{f_{c}(y_{i})}$
where $K$ is the number of inference, $f_{c}$ is the counting function.
Additionally, to assess whether the system’s output is intuitive and whether the output format meets user needs, we introduce a metric named format accuracy. Specifically, we aim for the output to be presented in JSON format, with clearly delineated key-value pairs. In this paper, the format should be {api_name: [Recommended API]}. Any alterations to this format, including missing symbols or changing the order of placement, are all considered as formatting errors. This structure is designed to facilitate user navigation, allowing for easy selection of desired sections.
6 Experiments
In this section, we conduct experiments on our collected dataset, evaluating the performance of our system MLLM-Tool. We start by introducing the experiment setup in Sec. 6.1 and data processing in Sec. 6.2. In Sec. 6.3, we assess MLLM-Tool’s ability on our benchmark, followed by results analysis in Sec. 6.4. And we conduct qualitative experiment and visualize the result in Sec. 6.5.
6.1 Experiment Setup
We list the setup, including some hyperparameters used to train our model in Table 4. The setup follows  [29] and adjusts based on our computation resources (4 $\times$ A40 40G GPUs). Especially, given the larger parameter size of the 13B model, to enhance experimental performance, we increase the epoch to 10.
@@ 6.1 Experiment Setup
6.2 Data processing
For multimodal inputs, we adopt ImageBind’s  [9] data processing approach. For video, we evenly split into five clips and randomly sample 2 frames from each clip to cover the full length, similar to other Vision Transformer (ViT)  [6] -based works  [32, 8] with video input. For audio, we sample each input audio at 16KHz. Subsequently, we capture a log mel spectrogram featuring 128 frequency bins, following  [10]. Then, the spectrogram can be seen as a 2D image. In this way, we can use ViT for these three multimodal encoders. A modality-specific linear projection head is added to each modality’s encoder to align the feature into a uniform 1024-sized dimension.

### v3

@@ 8.3 API Function Boundaries
One-to-many Situation Identification. Our approach, driven by GPT-4, originates from each API’s functional description and generates corresponding instructions that adhere strictly to the API’s functional boundaries. Hence, consideration is primarily given to other APIs with the same function in one-to-many scenarios.
8.4 Prompt Construction
Below, we list the detailed prompts used for generating instructions. The prompts fed to GPT-4 are composed of (1)the API call’s relevant information, including its name, a concise functional description, and training datasets if mentioned in the model card, (2)some sample instructions, and (3)additional requirements. The sample instructions are leveraged to foster format standardization and enhance the quality of GPT-4’s outputs. So, we invite some expert annotators to craft high-quality instructions in advance and randomly choose two each time as exemplary inputs to GPT-4. We design two sets of prompt templates in our dataset, mindful of prompt ambiguity cases. For APIs with multimodal inputs susceptible to prompt ambiguity, we strive to balance the final selection of ten instructions, maintaining a 1:1 ratio between unambiguous and ambiguous instructions; the corresponding prompt template is shown in Figure 8 and Figure 9 respectively. We emphasize that query ambiguity pairs are pairs with identical textual inputs but varied multimodal inputs.
8.5 Visualization of Data Distributions
We tally the instances of each ambiguity type in Figure 7(b), amassing a total of 2533 ambiguity pairs. Predominantly, these are concentrated within conditional cases involving image inputs, comprising 1530 pairs, since our dataset covers 11 distinct condition types of ControlNet  [42], and their various combinations, achieving a considerable quantity.
@@ 8.5 Visualization of Data Distributions
Particularly for the testing set, we depict in Figure 7(a) and Figure 7(c) the distribution of API quantities across each testing subset. These subsets are constructed based on three distinct partitioning criteria: ambiguity types, API option number, and modality. It is worth noting that the without ambiguity category consists of two parts: cases with purely textual inputs that inherently lack ambiguity (1,452 instances) and cases with non-textual inputs where ambiguity is absent (412 instances).
8.6 Crowdsourcing Details
In this study, we employ eight annotators with NLP or CV backgrounds, with an estimated hourly 7.10\$ compensation and an up to 20% bonus for those with high annotation quality. We provide a comprehensive training document and record a video to guide them through the process of annotation. We also instruct them to avoid using private information or license-risky data from unknown or sensitive sources. All annotators are presented with a consent form and must agree to join the project. We set up a cross-checking mechanism to correct errors. Moreover, our team members conduct manual checking and optimization. If the data is collected from a website, we restore the website links and upload the metadata when releasing the dataset. We claim we only use the data for academic research instead of commercial usage.
8.7 Collection of text-matched multimodal input
Our non-text data sources are from either the dataset API used or Google Images. We prioritize collecting from open-source datasets by manually checking that the data content is consistent with the instructions generated by GPT. For the latter, we can retrieve the data by keywords in the Google database, and for these data, we pay more attention to its license to ensure its legality.
@@ 8.7 Collection of text-matched multimodal input
9 Experiment Setup and Data Processing
We list the setup, including some hyperparameters used to train our model, in Table 9. We adopt ImageBind’s data processing approach for the multimodal inputs. We evenly split the video into five clips and randomly sampled two frames from each clip to cover the entire length, similar to other Vision Transformer (ViT)-based works with video input. For audio, we sample each input audio at 16KHz. Subsequently, we capture a log mel spectrogram featuring 128 frequency bins. Then, the spectrogram can be seen as a 2D image. In this way, we can use ViT for these three multimodal encoders. A modality-specific linear projection head is added to each modality’s encoder to align the feature into a uniform 1024-sized dimension.
10 Additional Experimental Results
10.1 Results of Multiple Options Categorized by Input Modality
Regarding the condition of whether multiple API options exist, we further refine the testing subsets—originally divided based on the presence or absence of multiple API options—by introducing an additional criterion based on input modality to gain deeper insights into the issue. We select two versions of the Vicuna model with different parameter sizes, and the corresponding performance is presented in Table 10. From the results, we observe that for the case where multiple API options exist, the performance is consistently better than the case with only one API option, regardless of the input modality. This validates the argument presented in the main text. On one hand, the presence of multiple API options reduces the problem’s complexity and increases the likelihood of selecting the correct API. On the other hand, during training, the data involving multiple API options benefits from being split into multiple one-to-one training strategies, further improving performance.
Model API Options Input Modality Acc
Vicuna-7B one-to-one Video 55.56
Vicuna-7B one-to-many Video 83.33
Vicuna-7B one-to-one Audio 78.57
Vicuna-7B one-to-many Audio 96.32
Vicuna-7B one-to-one Image 75.82
Vicuna-7B one-to-many Image 96.30
Vicuna-7B one-to-one Text 58.24
Vicuna-7B one-to-many Text 86.43
Vicuna-13B one-to-one Video 44.44
Vicuna-13B one-to-many Video 91.67
Vicuna-13B one-to-one Audio 75.00
Vicuna-13B one-to-many Audio 95.59
Vicuna-13B one-to-one Image 86.81
Vicuna-13B one-to-many Image 96.10
Vicuna-13B one-to-one Text 70.21
Vicuna-13B one-to-many Text 90.24
Table 10: The performance of the testing subsets based on the joint criterion of the presence of multiple API options and the input modality.